OpenStudy (anonymous):

i need 25 minutes of fast help. OS people i am counting on you for help

OpenStudy (anonymous):

How is the ultimate goal of the Hindu achieved? Brahma is achieved through joining with dharma. Moksha is achieved through obeying karma. Brahma is achieved through good moral acts. Moksha is achieved through joining with Brahma. Which of these houses the largest migration of mammals in the world? the Great Rift Valley the Sahara Mount Kilimanjaro the Serengeti Why might the Greek philosophy of Stoicism have been popular with Roman citizens? It called for harsh penalties for those who displayed anger or greed. It suggested that the public revolt against their leaders. It focused on characters who transformed greatly. It expressed the belief that all people should be equal.

OpenStudy (anonymous):

What was the main reason for the Kingdom of Axum's success and growth? It had the greatest army of any kingdom at the time. It was located along major trade routes. It had gold and silver mines in its territory. It made a strong alliance with the Roman Empire. Which of these led to the growth of the Roman Empire? Roman rule brought wealth to their conquered lands and people. Roman rule allowed conquered peoples to learn about democracy. The Romans did not tax the people living in the lands they conquered. The Romans were tolerant of most religions of conquered peoples. How did the rule of Asoka affect the empire? He stopped warfare and had Buddhist shrines built throughout the Mauryan Empire. He produced a legal code that governed all people who lived in the Mauryan Empire. He ordered the killings of thousands of people in lands outside the Mauryan Empire. He imposed strict rules and punishments on people throughout the Mauryan Empire.

OpenStudy (anonymous):

i have more would you like to help or not?

OpenStudy (anonymous):

@jagr2713 @King.Void. @nincompoop @TheSmartOne

TheSmartOne (thesmartone):

Is this a test? :)

OpenStudy (anonymous):

a review test and i need help

TheSmartOne (thesmartone):

What do you mean a review test? o_0

OpenStudy (anonymous):

it is a review test before you take the actual test. a lot of people on here ask for help on these test.

TheSmartOne (thesmartone):

I found the answers to this test, I complied them here, enjoy! :D http://pages.citebite.com/a4o3b6r4a2tjy

OpenStudy (anonymous):

wait what? what did i do

OpenStudy (anonymous):

you banned me didn't you thx what about the 60,000 people that do it all the time huh

TheSmartOne (thesmartone):

I don't have any powers to ban anyone.

OpenStudy (anonymous):

what is this are you playing the rick roll song or is this a warning tell me why you sent me there

TheSmartOne (thesmartone):

To tell you that you can't cheat on tests. And if you see anyone else cheating on their tests, report them c:

OpenStudy (anonymous):

i am pretty sure that is what everyone does is take a question from their test and type it on openstudy. everyone does it

OpenStudy (anonymous):

the only time i get someone on a quiz to finally talk is for me to get a warning. not fair

OpenStudy (greencat):

What school asks these questions?

OpenStudy (anonymous):

flvs

OpenStudy (greencat):

Thank you. Now, we can verify the proper consequences for this form of violation. Please abstain from this behavior in the future.

OpenStudy (greencat):

Checkmate.

OpenStudy (anonymous):

it is a review test

OpenStudy (anonymous):

you know what this is so unfair how about you report everyone in OS because in one point you cheated in your life too

OpenStudy (greencat):

Rule #1 of Cheating: Don't let anyone know you are cheating.

OpenStudy (anonymous):

i am not cheating

OpenStudy (anonymous):

it doesn't matter what grade i get on a review test

OpenStudy (greencat):

If the grade doesn't matter, why cheat?

OpenStudy (anonymous):

this is for social science and i am failing to badly in this class and no one in openstudy seems to care i lose

OpenStudy (anonymous):

*if i lose

OpenStudy (anonymous):

@GreenCat

OpenStudy (greencat):

You lose.

OpenStudy (anonymous):

i failed my review and final exam thanks to you. you gave me no hope to pass, did i mention i cried last night.

OpenStudy (greencat):

You did not mention that.

OpenStudy (causticsyndicalist):

lml

OpenStudy (anonymous):

why is everyone giving the cashmere cat medals?

OpenStudy (causticsyndicalist):

Da wat? Dude, he's obviously a green cat.

OpenStudy (anonymous):

i never asked for your help and next time you type something that hurts my emotions i am reporting you

OpenStudy (sleepyjess):

\(\color{blue}{\text{Originally Posted by}}\) @musicmaster i failed my review \(\color{blue}{\text{End of Quote}}\) Reviews are for you to do on your own to see your weaknesses. \(\color{blue}{\text{Originally Posted by}}\) @musicmaster and final exam thanks to you. you gave me no hope to pass, did i mention i cried last night. \(\color{blue}{\text{End of Quote}}\) We are not allowed to help on exams, it is against OpenStudy's Anti-Cheating Policy

OpenStudy (anonymous):

i never asked for help on my exam just the review and no i can still fail my reviews it is not a pretest

OpenStudy (sleepyjess):

We are not allowed to help on reviews

TheSmartOne (thesmartone):

Also, the thing is, you only wanted the answers. You didn't come here for leraning. :p

TheSmartOne (thesmartone):

learning*

OpenStudy (anonymous):

i didn't say that i said none of those so either you assumed that i only wanted answers or you were going to give me the answers but not telling me how to work it out. any of you good in math go ahead teach me please. i am not posting the questions but tell me about math

TheSmartOne (thesmartone):

Math is a huge subject. What area in math do you need help in?

OpenStudy (sleepyjess):

How to find the slope of a line segment: Finding the slope of a line segment given 2 points is quite simple using the formula \(m = \dfrac{y_2-y_1}{x_2-x_1}\). Let's use the points (5, 2) and (8, 9). (5, 2) is \((x_1, y_1)\) and (8, 9) is \((x_2, y_2)\). Now we use those and substitute the values into the formula. Now we get \(m = \dfrac{9-2}{8-5}\). Now just simplify the numerator and denominator \(m = \dfrac{7}3\)

OpenStudy (greencat):

\[\cos(\alpha+\beta)=\cos \alpha \cos \beta - \sin \alpha \sin \beta\]

OpenStudy (sleepyjess):

1+2=? If you have 1 chocolate in 1 hand, and 2 chocolates in the other hand, how many chocolates do you have? 3!

OpenStudy (greencat):

\[\cos(\alpha-\beta)=\cos \alpha \cos \beta + \sin \alpha \sin \beta\]

OpenStudy (greencat):

x+x=2 2x=2 x=1

OpenStudy (anonymous):

sleepyjess and greencat are now reported for being rude and hurting my feelings

OpenStudy (greencat):

;-; You're the one who was trying to cheat.

OpenStudy (greencat):

There are flvs people over here that could take action for you cheating. I was watching out for you, bud.

OpenStudy (sleepyjess):

You said to teach you math...

TheSmartOne (thesmartone):

Report me to while you're at it, you're only getting a moderator to come to your post and find out that you tried to cheat c:

OpenStudy (sleepyjess):

I guess my slope explanation wasn't good enough ;-;

OpenStudy (anonymous):

again i was not cheating. name once that you were trying to help me. the checkmate part then you were mocking me then you were being a smart aleck and now you trying to play kiss up

OpenStudy (anonymous):

wow i sure wonder what 1+2 is hmm lets make sure that SMART ALEK GETS REPORTED

OpenStudy (sleepyjess):

I have no idea what level math you're at, so how was I supposed to know? You just said math

TheSmartOne (thesmartone):

^ ikr? We start trying to explain to you calc and you'll be like :insert confused face:

OpenStudy (anonymous):

yeah i knew what it meant i am 7th grade is it that hard. you think a 2 year old would get online and type like a 7th grader

OpenStudy (greencat):

The "checkmate part" was me trying to act cool to forget how lame i am...

OpenStudy (greencat):

;-;

OpenStudy (anonymous):

well it just made you a jock the humble avenger

OpenStudy (greencat):

hmm You are in 7th grade. You are probably not even old enough to be on OS.

TheSmartOne (thesmartone):

You have a lot to learn young grasshopper.

OpenStudy (anonymous):

i am 13

OpenStudy (greencat):

O_o Ok.. I'll take your word.

OpenStudy (anonymous):

last time i saw the Coc it said nothing about having your age

OpenStudy (anonymous):

i was born in 2002 on February 14

TheSmartOne (thesmartone):

The Terms and Conditions has the age. That is what you agreed to when you signed up.

OpenStudy (greencat):

http://prntscr.com/77e55w

OpenStudy (kidrah69):

@e.mccormick

OpenStudy (kidrah69):

End of post.

OpenStudy (anonymous):

ok i am done i want anymore ambassadors

OpenStudy (anonymous):

I SAID NO MORE!!!

OpenStudy (causticsyndicalist):

Looks like I'm still up for torturing you then. ;-)

OpenStudy (anonymous):

not now caustic

OpenStudy (kidrah69):

Please stop spamming the post. Leave it.

OpenStudy (anonymous):

i cant see your post cause you let another moderator in

eclipsedstar (eclipsedstar):

\(\color{blue}{\text{Originally Posted by}}\) @sleepyjess We are not allowed to help on reviews \(\color{blue}{\text{End of Quote}}\) I never knew that.

OpenStudy (kidrah69):

Neither did i .-.

OpenStudy (anonymous):

hmm TSO is a moderator and he is putting the coc aside

TheSmartOne (thesmartone):

Yes, I'm a moderator. c;

OpenStudy (kidrah69):

Ambassadors NOT mod

eclipsedstar (eclipsedstar):

TSO is not a mod.

OpenStudy (anonymous):

same thing they all boss people around anyways

OpenStudy (kidrah69):

Have a nice day now stop spamming.

OpenStudy (anonymous):

the subject is closed it probably below 600 questions by now no one will ever see it. so in my book this is okay to do. what it can't do is be opened with a mass tagging

OpenStudy (anonymous):

yeah kidrahs gone!!!!

TheSmartOne (thesmartone):

You know the more you reply to your own post, she'll get a notification and come back? lel

eclipsedstar (eclipsedstar):

I see a fault here. Why is everyone ganging up on one user?

TheSmartOne (thesmartone):

Define everyone

eclipsedstar (eclipsedstar):

TSO, etc etc.

TheSmartOne (thesmartone):

So I'm everyone? :p

eclipsedstar (eclipsedstar):

The user already said that he was taking a review. It is not a test. Why was the CoC brought up anyways?

TheSmartOne (thesmartone):

CoC was never brought up fyi